OpenStudy (anonymous):

please help. Find the solutions of the equation. 1/2x 2 - x + 5 = 0

OpenStudy (owlcoffee):

like this?: \[\frac{ 1 }{ 2x }2-x+5=0 \]

OpenStudy (anonymous):

1/2x^2-x+5=0 there's an exponent.

OpenStudy (khalilforthewin):

hello

OpenStudy (owlcoffee):

\[(\frac{ 1 }{ 2x })^{2}-x+5=0\] Like that? or is the rest also an exponent?

OpenStudy (khalilforthewin):

ik the answer

OpenStudy (tkhunny):

\(\dfrac{1}{2}x^{2} - x + 5 = 0\) Maybe multiply by 2? \(x^{2} - 2x + 10 = 0\)

OpenStudy (khalilforthewin):

that

OpenStudy (anonymous):

@Owlcoffee thats right! & @khalilforthewin can u explain?

OpenStudy (owlcoffee):

okay, let's see: \[(\frac{ 1 }{ 2 })x^2-x+5=0\] This can be solved using Tkhunny's form: by multiplying the whole thing by 2: \[x^2-2x+10=0\] And then, apllying the quadratic formula: \[x=\frac{ 2 \pm \sqrt{2^2-4(1)(10)} }{ 2 }\]

OpenStudy (anonymous):

x=1±3i

OpenStudy (owlcoffee):

good! it's 1+/- 3i well done!

OpenStudy (anonymous):

@Owlcoffee yay! thank you! (:

OpenStudy (khalilforthewin):

back off she is mine

OpenStudy (owlcoffee):

Girls are no trophies :(

OpenStudy (khalilforthewin):

no....no i didnt mean it like that

OpenStudy (anonymous):

@khalilforthewin @Owlcoffee lol (: